# Fidelity Sweet Spot in Transmon Qubit Rings under Strong Connectivity Noise

### v1

@@ Abstract
We investigate the fidelity of quantum operations in transmon qubit systems, focusing on both SWAP and general gate operations. Our results reveal a distinct fidelity sweet spot that emerges even under strong noise, indicating that optimal circuit depth can enhance gate performance. We further demonstrate that specific initial states, particularly those with favorable symmetry or entanglement structure, yield higher fidelity, reaching levels compatible with quantum error-correction thresholds. Finally, we introduce a supervised machine-learning framework capable of predicting the positions of fidelity sweet spots, enabling efficient optimization of circuit durations across different device configurations.
@@ I Introduction
Superconducting qubits are a leading platform for quantum computation owing to their precise controllability and the potential for long coherence times under optimized conditions Blais et al. (2004, 2007); Schuster et al. (2007). They rely on Josephson junctions to enable fast electrical manipulation of quantum states, forming the foundation for scalable gate-based architectures Josephson (1962, 1974); You and Nori (2011). Among these implementations, the transmon qubit—a refined version of the Cooper-pair box—has become the dominant design because of its reduced charge sensitivity and enhanced coherence Koch et al. (2007); Paik et al. (2011). Typically fabricated from aluminum and operated at temperatures below 30 mK, transmon circuits achieve strong coupling to coplanar waveguide (CPW) resonators by concentrating the electric field near the qubit Yoshihara et al. (2017); Kollár et al. (2019). This configuration strengthens the dispersive interactions essential for quantum information processing Jaynes and Cummings (1963); Cummings (2013), but also introduces connectivity-related noise that can degrade fidelity during multi-qubit operations Throckmorton and Das Sarma (2022); Fu et al. (2024).
A critical challenge arises from the noise-fidelity trade-off: while reduction of noises favors shallower circuit depths, such shallow circuits inherently lack the computational capacity required for complex quantum algorithms. As an example, for SWAP gates–a key primitive for state transfer and manipulation Throckmorton and Das Sarma (2020); Foulk et al. (2022)–the gate duration scales as $\tau=\frac{\pi}{4J}$ (assuming $\hbar=1$), where $J$ quantifies qubit-qubit coupling. We use $\lambda_{0}$ to denote a typical noise amplitude, and one expects that for sufficiently large $J/\lambda_{0}$, the infidelity of quantum operations should decrease monotonically with increasing $J/\lambda_{0}$ Throckmorton and Das Sarma (2020); Foulk et al. (2022). For transmon devices, the threshold is around $J/\lambda_{0}\gtrsim 100$ Barends et al. (2014); Mills et al. (2022); Noiri et al. (2022). However, sustaining this ratio in practical transmon devices remains cumbersome in strong coupling regimes Caldwell et al. (2018); Kong et al. (2015).
Hence in the noisy intermediate-scale quantum era, identifying operational sweet spots, particularly in the regime $10\lesssim J/\lambda_{0}\lesssim 100$, is crucial. For a fixed $J$, this means the system endures stronger noise than in the regime $J/\lambda_{0}\gtrsim 100$, yet it also permits a longer total duration for gate execution, thereby facilitating the realization of quantum algorithms if a sweet spot exists.
In this work, we demonstrate that a fidelity sweet spot in the intermediate-coupling regime $10\lesssim J/\lambda_{0}\lesssim 100$ emerges in fully connected transmon rings with varying qubit numbers Deng et al. (2015); Kong et al. (2015); Zhu et al. (2017). We begin by analyzing a sequence of SWAP gate operations and then extend the study to more general, randomly generated manipulations Throckmorton and Das Sarma (2020); Foulk et al. (2022). For any fixed system and initial state, such a sweet spot in $J/\lambda_{0}$ consistently appears; although its location may shift slightly between different operations, the displacement is negligible. Further analysis of fidelity across diverse initial states shows that, although the sweet spot’s position may shift, its associated infidelity for certain states becomes exceptionally low—approaching the error-correction threshold of 99.9% Kandel et al. (2019); Sigillito et al. (2019). These insights suggest that there exists an optimal time duration of quantum circuits (circuit duration or depth) executed in such systems, and could steer the design of future circuits toward greater efficiency and accuracy under a wide range of operating conditions.
Recognizing the experimental importance of the location of such sweet spots under variations in CPW coupling strength and noise distribution, we apply a supervised machine-learning strategy to predict the fidelity sweet spot for a given quantum device using only a small amount of simulation data Czarnik et al. (2021); Strikis et al. (2021); Cincio et al. (2021). This approach supplies the optimal circuit duration for new parameter regimes without the heavy resource cost of exhaustive simulations or measurements for every configuration Zlokapa and Gheorghiu (2020); Rodrigues (2023).
The rest of the paper is organized as follows. In Sec. II, we introduce our model and methods. Next, in Sec. III.1, we present the results for fidelity sweet spots under SWAP gates. Then, we show the results for a more general case, i.e. randomly generated general operations in Sec. III.2. After that, we demonstrate the impact of different initial states on fidelity in Sec. III.3. A discussion about the mechanism for the sweet spot is presented in Sec. III.4. We then use a supervised neural network machine learning technique to predict the sweet spot and obtain solid results in Sec. IV. Finally, we conclude in Sec. V.
@@ II Model and method
Figure 1: Schematic of transmon-qubit devices with different numbers of qubits $L$. Each node represents a transmon qubit, with solid lines denoting cavity couplings and wavy lines denoting CPW couplings. Panel (a) highlights examples of a nearest-neighbor cavity coupling $J_{12}$ and a remote CPW coupling $K_{13}$.
Transmon qubits can be coupled in two primary ways: direct cavity coupling or CPW coupling Xiang et al. (2013); Koch et al. (2007); Majer et al. (2007); DiCarlo et al. (2009); Ristè et al. (2015). When coupled through CPW, the qubits may be spatially separated; under suitable conditions, the CPW can mediate strong indirect interactions, enabling long-range entanglement and remote transfer of quantum information Kollár et al. (2019). In this context, we regard the effective connectivity of neighboring qubits as arising from cavity coupling, whereas physically distant qubits interact via CPW coupling. We study three transmon-qubit devices with different number of qubits $L$, as shown in Fig. 1. Every vertex denotes a transmon qubit, with solid lines indicating direct cavity couplings and wavy lines indicating CPW-mediated couplings. Across these devices, adjacent qubits form a ring through cavity connections, while CPW links supply the long-range interactions essential for distributing entanglement. The control Hamiltonian is given by
@@
where $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}$ are Pauli matrices for qubit $i$, $\sigma_{i}^{+}=(\sigma_{x}+i\sigma_{y})/2$, $\sigma_{i}^{-}=(\sigma_{x}-i\sigma_{y})/2$. $\Omega_{i}^{x}(t)$ is single-qubit driving term, $\Delta_{i}(t)$ is the detuning, $\langle i,j\rangle$ denotes the adjacent qubits via cavity coupling $J_{ij}(t)$, while $\{i,j\}$ denotes the non-adjacent qubits via CPW coupling $K_{ij}(t)$ [cf. Fig. 1(a))]. Appendix A provides a derivation of Eq. (1) from the circuit Hamiltonian of transmon qubits.
Eq. (1) can generate any unitary $R\in\text{SU}(2^{L})$ Lloyd (1995); Dawson and Nielsen (2006), thus suffices for universal quantum computation. To implement a SWAP gate, we set $\Omega_{i}^{x}(t)=\Delta_{i}(t)=K_{ij}\equiv 0$, and $J_{ij}=J$ for the targeting adjacent $\langle i,j\rangle$ pairs being swapped. For a general operation $R$ which is randomly chosen, the parameters $\Omega_{i}^{x}(t)$, $\phi_{i}(t)$, $\Delta_{i}(t)$, $J_{ij}(t)$ and $K_{ij}(t)$ are computed according to the procedure outlined in Appendix B.
Figure 2: Schematic of a two-qubit quantum circuit subject to noise. The circuit consists of a Hadamard gate followed by a CNOT gate, both operating under noisy conditions. In this study, the noise primarily represents connectivity-induced fluctuations in transmon-qubit systems.
We now examine how connectivity-induced noise, most notably parasitic capacitance noise in cavity-coupled pairs and microwave-photon loss in CPW channels Kollár et al. (2019); Kong et al. (2015), impacts quantum operations. Although many noise sources exist, our work isolates the effects arising specifically from connectivity. For a circuit meant to implement the unitary $R$, noisy processes (represented by $U^{\dagger}$) transform the evolution into $U^{\dagger}R$, altering the final state. Figure 2 shows a two-qubit circuit illustrating this effect (shaded regions). The circuit performs a Hadamard gate followed by a CNOT gate, and both gates are subject to noise. The corresponding effective noise model is:
@@
The dominant source of noise between qubits are realized by parasitic capacitance ($\lambda_{ij}^{J}(t)$) and photon loss ($\lambda_{ij}^{K}(t)$) (cf. Appendix A), which are independent from control operations Johnson et al. (2011); Yan et al. (2018); Zhang et al. (2019); Zhao et al. (2025a); Hai et al. (2025). The $\lambda_{ij}^{J}$ term originates primarily from parasitic capacitance within the device, with its amplitude scaling as $\lambda_{ij}^{J}\propto C_{g}^{\text{para}}$, where $C_{g}^{\text{para}}$ represents the parasitic capacitance between qubits $i$ and $j$ Zhao et al. (2025b). The CPW noise term $\lambda_{ij}^{K}$ stems from photon loss through the superconducting waveguides, following the fluctuation amplitude $\lambda^{K}_{ij}\propto Q^{-1/2}$, with $Q$ being the quality factor Zhang et al. (2019). Due to fabrication tolerances across processor batches Kong (2018), the characteristic ratio $\mathcal{R}=\lambda_{ij}^{K}/\lambda_{ij}^{J}$ ranges between 3 and 20. We accordingly maintain this ratio in our numerical simulations. The frequency shift $\delta\omega_{i}$ predominantly arises from charge and flux noise, with typical values relative to CPW coupling $\frac{\delta\omega_{i}}{\lambda_{ij}^{K}}\sim 0.1\%$. This ensures that $\delta\omega_{i}$ acts as a minor perturbation compared to $\lambda_{ij}^{K}$.
In transmon systems, typical experimental noises are much slower than the characteristic timescale of the quantum operations Koch et al. (2007); Kong (2018); Wang et al. (2022); Throckmorton and Das Sarma (2022). Therefore, we primarily focus on the effect of quasi-static noise, which is characterized by a normal distribution. Specifically, we choose $\lambda_{ij}^{K}(t)\sim\mathcal{N}(\lambda_{ij,0}^{K},\sigma_{K}^{2})$ and $\lambda_{ij}^{J}(t)\sim\mathcal{N}(\lambda_{ij,0}^{J},\sigma_{J}^{2})$ Fu et al. (2024).
We define the fidelity $F$ as Throckmorton and Das Sarma (2020)
@@
where $|\Psi_{0}\rangle$ is the initial state, $U^{\dagger}$ is due to the presence of noise, and $R$ is the corresponding operation we consider. We randomly generate 1,500 samples of the noise values from the Gaussian distributions and then average them to obtain the mean fidelity $\overline{F}$. We take the energy unit as $\lambda_{0}$, and take $\lambda_{ij,0}^{K}=\lambda_{0},\lambda_{ij,0}^{J}=0.1\lambda_{0},\delta\omega_{i}=10^{-4}\lambda_{0}$, with $\sigma_{K}=0.01\lambda_{ij,0}^{K},\sigma_{J}=0.01\lambda_{ij,0}^{J}$. These parameter values are consistent with experiments for transmon devices, as detailed in Appendix A. In Sec. III.1, we consider the case of executing a series of SWAP gates Throckmorton and Das Sarma (2020). We then extend our study to the case of general operations in Sec. III.2.
Note that we assume all the quantum operations are performed perfectly (i.e. no noise terms in Eq. (1)). As far as the initial states are concerned, we select several commonly used states both experimentally and theoretically, such as $\left|\Psi\right\rangle_{1}=\left|\uparrow\downarrow\downarrow\cdots\right\rangle$, $\left|\Psi\right\rangle_{\mathrm{S}}=\left|S\right\rangle\left|\downarrow\downarrow\cdots\right\rangle$ where the first two qubits are in a singlet state $\left|S\right\rangle=\frac{1}{\sqrt{2}}(\left|\uparrow\downarrow\right\rangle-\left|\downarrow\uparrow\right\rangle)$ and the rest L2L-2 qubits are spin down, and $\left|\Psi\right\rangle_{\mathrm{GHZ}}=\frac{1}{2}(\left|\uparrow\uparrow\cdots\right\rangle+\left|\downarrow\downarrow\cdots\right\rangle)$.
@@ III Results
III.1 Sequence of SWAP operations

### v2

# Enhancing Circuit Fidelity in Transmon Qubit Rings via Operation Duration Tuning under Strong Connectivity Noise

@@ Abstract
Superconducting transmon qubits are a promising platform for quantum computation, yet they face significant fidelity degradation due to connectivity noise, particularly in the intermediate coupling regime where noise levels are substantial. While prior works suggest that high fidelity requires operating in regimes with strongly suppressed noise, maintaining such conditions under practical experimental constraints remains challenging. To address this, we investigate quantum gate operations in fully connected transmon rings, examining both SWAP and general circuits. Our study reveals that fidelity can be significantly enhanced by tuning gate operation durations, with local maxima emerging even under strong noise conditions. These fidelity enhancements occur consistently across different qubit numbers and operation types, and for specific initial states—particularly those with favorable symmetry or entanglement properties—the achieved fidelities approach quantum error correction thresholds. Furthermore, we develop a supervised machine learning model that accurately predicts the optimal operation durations for new devices, enabling efficient optimization without extensive experimental simulations. These results provide a pathway toward robust quantum circuit design in noisy experimental environments.
Keywords: transmon qubit, quantum circuit fidelity, connectivity noise, operation duration tuning, machine learning, superconducting circuits
@@ I Introduction
Superconducting qubits represent a leading platform for scalable quantum computation, combining precise controllability with the potential for extended coherence under optimized conditions Blais et al. (2004, 2007); Schuster et al. (2007). Building on this foundation, these systems exploit Josephson junctions to enable rapid electrical manipulation of quantum states, forming the basis for gate-based architectures Josephson (1962, 1974); You and Nori (2011). Among various implementations, the transmon qubit—a refined Cooper pair box design—has emerged as the dominant architecture due to its reduced charge noise sensitivity and enhanced coherence properties Koch et al. (2007); Paik et al. (2011). Typically fabricated from aluminum and operated below 30 mK, transmons achieve strong coupling to coplanar waveguide (CPW) resonators by concentrating electric fields near the qubit Yoshihara et al. (2017); Kollár et al. (2019). While this configuration enhances dispersive interactions vital for quantum information processing Jaynes and Cummings (1963); Cummings (2013), it simultaneously introduces connectivity induced noise that significantly degrades fidelity in multiqubit operations Throckmorton and Das Sarma (2022); Fu et al. (2024).
This inherent trade-off between connectivity and noise leads to a critical challenge: although shorter gate durations minimize noise accumulation, they fundamentally limit the complexity of implementable quantum operations Throckmorton and Das Sarma (2020); Foulk et al. (2022). To quantify this trade-off, consider the SWAP gate—a key primitive for state transfer—whose duration scales as $\tau=\pi/(4J)$ (with $\hbar=1$), where $J$ represents the qubit-qubit coupling strength. For a fixed noise amplitude $\lambda_{0}$, conventional wisdom suggests that infidelity should decrease monotonically with increasing $J/\lambda_{0}$, with high-fidelity operation requiring $J/\lambda_{0}\gtrsim 100$ in transmon devices Barends et al. (2014); Mills et al. (2022); Noiri et al. (2022). However, maintaining such high ratios proves experimentally challenging in practical strong-coupling regimes Caldwell et al. (2018); Kong et al. (2015), particularly in the noisy intermediate-scale quantum era where precise control over environmental parameters is limited.
To address this fundamental limitation, we demonstrate that optimal operation points emerge in the more accessible intermediate-coupling regime ($10\lesssim J/\lambda_{0}\lesssim 100$) for fully connected transmon rings Deng et al. (2015); Kong et al. (2015); Zhu et al. (2017). We define an optimal operation point as a specific gate duration (or equivalent $J/\lambda_{0}$ ratio) where gate fidelity reaches a local maximum despite substantial noise, representing a dynamic balance between operation time and decoherence. Crucially, this approach contrasts with the conventional strategy of simply minimizing gate duration and enables longer duration operations under realistic noise conditions. Our investigation begins with SWAP gate sequences and extends to general randomly generated operations Throckmorton and Das Sarma (2020); Foulk et al. (2022), revealing that optimal points appear consistently across different systems with positions largely independent of the specific circuit implementation. Remarkably, for certain initial states—particularly GHZ-type entangled states—fidelity at these optimal points approaches the error-correction threshold of 99.9% Kandel et al. (2019); Sigillito et al. (2019), highlighting the important role of state-circuit symmetry in noise resilience.
While these optimal operation points offer significant advantages, their experimental implementation faces practical challenges due to device-to-device variations in CPW coupling strengths and noise distributions. To overcome this issue, we integrate a supervised machine learning approach that accurately predicts optimal operation points using limited simulation data Czarnik et al. (2021); Strikis et al. (2021); Cincio et al. (2021). This methodology avoids the need for exhaustive simulations for each parameter configuration Zlokapa and Gheorghiu (2020); Rodrigues (2023) and facilitates adaptive optimization of gate durations across diverse experimental setups.
The remainder of this paper is organized as follows: Section II details the model and method. Sections III.1 and III.2 present comprehensive results for SWAP gates and general operations, respectively. Section III.3 explores the dependence of fidelity on initial states, Section III.4 discusses the underlying physical mechanisms, and Section IV describes the machine learning framework for optimal point prediction. Finally, Section V summarizes our key findings and discusses their implications for future quantum processor design.
@@ II Model and method
Figure 1: Schematic of transmon qubit devices with different numbers of qubits $L$. Each node represents a transmon qubit, with solid lines denoting nearest-neighbor (cavity-mediated) couplings and wavy lines denoting long-range (CPW-bus-mediated) couplings. Panel (a) highlights examples of a nearest-neighbor coupling $J_{12}$ and a remote CPW-bus coupling $K_{13}$.
Transmon qubits are coupled through two primary mechanisms: nearest-neighbor couplings mediated by on-chip superconducting resonators and long-range couplings mediated by CPW transmission lines acting as quantum buses. In experimental devices, nearest-neighbor couplings are often implemented using dedicated tunable couplers (which are themselves transmon qubits) for compactness and fast control. In our theoretical model, we abstract this implementation detail and represent the nearest-neighbor exchange interaction as a cavity-mediated coupling with strength $J_{ij}(t)$. The long-range CPW-mediated couplings, represented by $K_{ij}(t)$, enable strong indirect interaction between spatially separated qubits and have been used in quantum bus architectures to facilitate long-range entanglement and quantum information transfer Kollár et al. (2019). While CPW transmission lines are physically larger than qubits, they provide a viable route to implement all-to-all connectivity in modular quantum processors. Our study focuses on the theoretical implications of such hybrid connectivity and its associated noise on circuit fidelity.
We study three transmon qubit devices with varying qubit numbers $L$, as shown in Fig. 1. Each vertex represents a transmon qubit, with solid lines indicating nearest-neighbor couplings and wavy lines indicating long-range CPW-bus-mediated couplings. The devices form ring structures where adjacent qubits are connected via nearest-neighbor couplings, while CPW links provide essential long-range interactions for quantum information distribution.
The control Hamiltonian governing these systems is:
@@
where $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}$ are Pauli matrices for qubit $i$, $\sigma_{i}^{\pm}$ are raising/lowering operators, $\Omega_{i}^{x}(t)$ is the single qubit drive amplitude, $\Delta_{i}(t)$ is the detuning, $\langle i,j\rangle$ denotes adjacent qubits coupled via a resonator or tunable coupler with strength $J_{ij}(t)$, and $\{i,j\}$ denotes non-adjacent qubits coupled via a CPW bus with strength $K_{ij}(t)$ [cf. Fig. 1(a)]. Appendix A provides the full derivation from circuit Hamiltonians.
We note that in our model, the coupling strengths $J_{ij}(t)$ and $K_{ij}(t)$ are treated as tunable control parameters. In actual devices, these couplings are often adjusted by changing the qubit frequencies (thus modifying the detuning from a shared resonator or bus), which introduces correlations between different couplings. However, our theoretical abstraction allows us to focus on the generic effects of connectivity noise. For SWAP gate sequences, we consider sequential operations where only one coupling is active at a time, which can be achieved by detuning non-target qubits, thereby decoupling the couplings in time. This is consistent with the commonly used ”activate/deactivate” control strategy in experiments. For general multi-qubit unitaries, the universal quantum computing theory ensures that any operation can be decomposed into single- and two-qubit gates, each of which can be realized by adjusting the couplings without fundamental limitations. Although correlations between couplings exist in actual devices, optimization algorithms that take these correlations into account can find pulse sequences to achieve the target operation. Our model, by assuming tunable control, captures the essential physics of connectivity noise and the existence of optimal operation points.
This Hamiltonian enables universal quantum computation by generating any unitary $R\in\text{SU}(2^{L})$ Lloyd (1995); Dawson and Nielsen (2006). For SWAP gates, we set $\Omega_{i}^{x}(t)=\Delta_{i}(t)=K_{ij}\equiv 0$ and $J_{ij}=J$ for the target qubit pairs. For general operations $R$, we compute the time-dependent parameters using the procedure in Appendix B.
Figure 2: Schematic of a two qubit quantum circuit subject to connectivity noise. The circuit (Hadamard followed by CNOT) operates under noisy conditions, with shaded regions indicating noise affected segments. In transmon systems, this noise primarily stems from connectivity induced fluctuations.
II.1 Connectivity noise model
The central challenge in achieving high fidelity operations arises from connectivity induced noise, which becomes particularly significant in the intermediate coupling regime ($10\lesssim J/\lambda_{0}\lesssim 100$). We focus on two dominant noise sources: parasitic capacitance noise in cavity coupled pairs and microwave photon loss in CPW channels Kollár et al. (2019); Kong et al. (2015). For a target unitary $R$, noisy processes transform the evolution to $U^{\dagger}R$, altering the final state as illustrated in Fig. 2.
The noise Hamiltonian is:
@@ II.1 Connectivity noise model
[…]
We note that in a more detailed physical model, tuning the qubit frequencies (via $\Delta_{i}(t)$ or via design parameters $\Delta_{ik}$) could modify the coupling strengths $g_{k}^{i}(\omega)$ and consequently the noise amplitudes. However, in our quasi-static noise model, we treat $\lambda_{ij}^{K}$ and $\lambda_{ij}^{J}$ as fixed parameters determined by static device properties. This simplification is valid when the dynamic frequency shifts used for gate operations are small compared to typical qubit-resonator detunings. Our study focuses on the effect of fixed noise amplitudes on circuit fidelity as a function of gate duration; a more detailed noise model that includes dependence on operational parameters is an important direction for future device-specific optimization.
Given the slow noise dynamics in transmon systems compared to gate operations Koch et al. (2007); Throckmorton and Das Sarma (2022), we model quasistatic noise using normal distributions: $\lambda_{ij}^{K}(t)\sim\mathcal{N}(\lambda_{ij,0}^{K},\sigma_{K}^{2})$ and $\lambda_{ij}^{J}(t)\sim\mathcal{N}(\lambda_{ij,0}^{J},\sigma_{J}^{2})$ Fu et al. (2024).
II.2 Fidelity metric and simulation parameters
We quantify operation quality using the state fidelity:
@@ II.2 Fidelity metric and simulation parameters
where $|\Psi_{0}\rangle$ is the initial state. For each parameter set, we generate 1,500 noise samples from Gaussian distributions and compute the mean fidelity $\overline{F}$.
Using $\lambda_{0}$ as the energy unit, we set $\lambda_{ij,0}^{K}=\lambda_{0}$, $\lambda_{ij,0}^{J}=0.1\lambda_{0}$, $\delta\omega_{i}=10^{-4}\lambda_{0}$, with relative fluctuations $\sigma_{K}=0.01\lambda_{ij,0}^{K}$ and $\sigma_{J}=0.01\lambda_{ij,0}^{J}$. These values align with experimental transmon devices (Appendix A). The key parameter $J/\lambda_{0}$—which directly relates to gate duration through $\tau\propto 1/J$—is varied across the intermediate coupling regime to identify optimal operation points.
We consider three classes of initial states:
  • Product state: $|\Psi\rangle_{1}=|\uparrow\downarrow\downarrow\cdots\rangle$
  • Entangled pair: $|\Psi\rangle_{\mathrm{S}}=|S\rangle|\downarrow\downarrow\cdots\rangle$ with $|S\rangle=\frac{1}{\sqrt{2}}(|\uparrow\downarrow\rangle-|\downarrow\uparrow\rangle)$
  • GHZ state: $|\Psi\rangle_{\mathrm{GHZ}}=\frac{1}{\sqrt{2}}(|\uparrow\uparrow\cdots\rangle+|\downarrow\downarrow\cdots\rangle)$
These represent common experimental preparations with varying entanglement structures, allowing us to probe state dependent fidelity behavior.
In Sec. III.1, we analyze SWAP gate sequences, while Sec. III.2 extends to randomly generated general operations. Throughout, we assume perfect control operations (no noise in Eq. (1)) to isolate connectivity noise effects.
@@ III Results
III.1 Optimal gate duration of SWAP operations